### v1

# RemoteReasoner: Towards Unifying Geospatial Reasoning Workflow

Abstract
Remote sensing imagery presents vast, inherently unstructured spatial data, demanding sophisticated reasoning to interpret complex user intents and contextual relationships beyond simple recognition tasks. In this paper, we aim to construct an Earth observation workflow to handle complex queries by reasoning about spatial context and user intent. As a reasoning workflow, it should be somewhat autonomous, where predefined ground-truth reasoning paths do not constrain the learning process. Furthermore, its architecture ought to be unified yet flexible, enabling the model to perform diverse reasoning tasks with distinct output formats through a single forward pass. Existing remote sensing approaches fail to address these requirements, as they rely on supervised fine-tuning paradigms that constrain the autonomy of reasoning. To this end, we propose RemoteReasoner, a flexible and robust workflow for remote sensing reasoning tasks. The design of RemoteReasoner integrates a multi-modal large language model (MLLM) for interpreting user instructions and localizing targets, together with task adaptation strategies that enable multi-granularity output generation. In contrast to existing methods, our framework is trained with reinforcement learning (RL) to endow the MLLM sufficient autonomy for precise reasoning. At the inference stage, our adaptation strategies enable diverse output formats at inference time without requiring task-specific decoders or further fine-tuning. Preliminary experiments demonstrated that RemoteReasoner achieves remarkable performance across multi-granularity reasoning tasks, including region-level and pixel-level. Additionally, our framework enables novel capabilities such as the contour extraction task beyond the reach of existing reasoning pipelines.
Keywords: Remote sensing, Reasoning workflow, Reinforcement learning
= Projects: Work in progress.
= Contact: yaoliang@hhu.edu.cn
1 Introduction
Remote sensing analysis has evolved beyond simple object detection Zou et al. (2023) or classification Adegun et al. (2023), demanding sophisticated reasoning capabilities to unlock its full potential for complex, real-world decision-making. Although traditional pixel- Yan et al. (2023) or object-centric Liu et al. (2025) approaches are effective for narrow tasks, they fundamentally lack the ability to interpret nuanced user intents, infer implicit spatial relationships, or dynamically chain logical steps to answer intricate queries. Reasoning Jaech et al. (2024), Xu et al. (2025) bridges this gap by enabling systems to dynamically decipher the underlying goals of ambiguous or high-level instructions (e.g., "Find areas at risk of flooding near critical infrastructure after heavy rain").
Numerous efforts Lai et al. (2024), Ren et al. (2024), Chen et al. (2024b) have explored reasoning tasks in the natural image domain. For instance, LISA Lai et al. (2024) utilizes multi-modal large language models (MLLMs) Carolan et al. (2024) for reasoning-driven segmentation tasks. However, these approaches rely on supervised fine-tuning (SFT) paradigms to train the models. Such frameworks constrain model autonomy, ultimately compromising generalization capabilities. This situation can be manifested in substantial performance degradation on out-of-distribution samples and catastrophic forgetting of general capabilities. Similarly, the remote sensing community has seen pioneering works Li et al. (2025), Yao et al. (2025b). SegEarth-R1 Li et al. (2025) introduces a geospatial pixel reasoning task that integrates a hierarchical MLLM and a mask decoder to generate both reasoning processes and segmentation outputs. However, it remains constrained by SFT paradigms and exclusively supports segmentation tasks.
These limitations motivate us to design a reasoning workflow that autonomously interprets user intent through self-directed thinking and efficiently generates geospatial interpretations in diverse output formats tailored for downstream intelligence analysis. Firstly, the requirement for self-thinking should be fulfilled through an open-ended learning paradigm. This aligns naturally with the emerging reinforcement learning (RL) framework Guo et al. (2025), Shen et al. (2025). Rather than SFT approaches that demand strictly annotated reasoning processes, RL explores the self-evolution potential of models to cultivate reasoning capabilities, thereby naturally enhancing model generalization. Secondly, to support flexible multi-task adaptation, we should avoid redundant training through task transformation at the decision-making level.
Followed by these insights, we propose RemoteReasoner, a reasoning workflow that integrates a robust MLLM and a flexible inference strategy. Building upon the pixel-level reasoning task proposed in SegEarth-R1 Li et al. (2025), RemoteReasoner further implements two novel tasks through autonomous reasoning: 1) region-level: geospatial region reasoning, and 2) object-level: geospatial contour reasoning. These three distinct tasks can be achieved through a single MLLM inference followed by a unified decision-level task transformation pipeline, significantly improving output efficiency for multi-granularity reasoning tasks.
During training, we utilize Group Relative Policy Optimization (GRPO) Shao et al. (2024) to fine-tune the reasoning model. Benefiting from our sophisticated reward function that balances output formats, localization accuracy, and quantity, RemoteReasoner demonstrates strong capabilities in autonomous reasoning and precise region-level localization. This training approach maintains the inherent generalization capability of MLLM, thereby endowing it with the ability to recognize out-of-distribution categories. During inference, we design a task transformation pipeline inspired by RemoteSAM Yao et al. (2025b), which enables the model to perform object- and pixel-level reasoning tasks, eliminating redundant computations across granularities.
Extensive experimental results demonstrate that RemoteReasoner achieves state-of-the-art performance on pixel-level and region-level reasoning tasks, with enhanced autonomous reasoning capabilities and superior generalization. For instance, RemoteReasoner achieves over 40% accuracy improvement in region-level geospatial reasoning tasks. Moreover, compared to existing methods, RemoteReasoner supports a broader range of tasks without introducing additional decoders.
The contributions are summarized as follows:
We propose RemoteResoner, a unified yet flexible reasoning workflow for Earth observation, which reveals that integrating pure RL and task transformation strategies could handle multi-granularity reasoning tasks.
We introduce two novel geospatial reasoning tasks, region reasoning and contour reasoning. We also provide the corresponding precisely annotated datasets to facilitate the advancement of remote sensing research.
Holistic evaluations demonstrate that RemoteReasoner achieves superior generalization capabilities and efficiency in handling multi-granularity reasoning tasks through a single forward pass.
2 Related Work
2.1 Reasoning in Large Language Models
Recent efforts Jaech et al. (2024), Guo et al. (2025), Ren et al. (2024), Lai et al. (2024), Li et al. (2025) have shown that reasoning is central to aligning large language models (LLMs) with complex instructions. Models like OpenAI-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) incorporate chain-of-thought(CoT) Wei et al. (2022) reasoning into their generation process, enabling flexible and interpretable multi-step inference. Both adopt reinforcement learning (RL) Kaelbling et al. (1996), Li (2017), Arulkumaran et al. (2017) to induce generalizable reasoning behaviors—OpenAI-o1 through RLHF and CoT monitoring, and DeepSeek-R1 via large-scale RL from scratch followed by multi-stage pretraining. These approaches exhibit strong reasoning capabilities but remain limited to natural language and vision tasks.
In remote sensing, models such as PixelLM Ren et al. (2024), LISA Lai et al. (2024), and SegEarth-R1 Li et al. (2025) extend reasoning to spatial domains. While effective for pixel- or region-level inference, they often support narrow task types and fixed granularity. Notably, SegEarth-R1 relies on supervised fine-tuning, making it less robust to distribution shifts and less adaptable in open-ended settings.
These limitations underscore the need for a more versatile geospatial framework. To this end, we propose RemoteReasoner, which supports unified and efficient multi-task, multi-granularity remote sensing understanding.
2.2 Remote Sensing Multi-modal Models
Recent remote sensing multimodal models Muhtar et al. (2024), Yao et al. (2025a), Zhan et al. (2024), Kuckreja et al. (2024), Yao et al. (2025b) demonstrate strong capabilities in handling diverse tasks across spatial granularities. Falcon Yao et al. (2025a) supports 14 vision-language tasks via multi-resolution encoding; GeoChat Kuckreja et al. (2024) and LHRS-Bot Muhtar et al. (2024) incorporate spatial prompts and geographic priors for improved region-level grounding; RemoteSAM Yao et al. (2025b) adopts a unified encoder for dense pixel-level segmentation; SkyEyeGPT Zhan et al. (2024) reformulates task input/output as natural language for broad task generalization. While effective in geospatial understanding, these models lack robust reasoning capabilities, limiting performance under ambiguous or compositional prompts.
While existing models demonstrate strong perception capabilities, they lack the reasoning ability to interpret complex prompts or adapt across granularities. RemoteReasoner addresses this gap by incorporating autonomous reasoning, enabling precise execution of complex natural language instructions and strong generalization across multi-granularity geospatial tasks.
3 RemoteReasoner
In this section, we introduce our method of building a workflow for multi-level geospatial reasoning tasks. The overall framework is represented in Fig. 1.
3.1 Data
The EarthReason Li et al. (2025) dataset is proposed alongside SegEarth-R1, containing 6 questions per sample with an average length of 20.86 words. It comprises 2,371 training, 1,135 validation, and 1,928 testing samples. However, this dataset is limited to pixel tasks. To address this, we generate detection and contour annotations through Mask2Bbox (M2B) Liu et al. (2024a) and Mask2Contour (M2C) Miao et al. (2025) strategies, respectively. Additionally, we design prompt templates combined with original questions and generated bounding boxes to form training samples. The segmentation and contour annotations are exclusively reserved for downstream testing evaluation.
3.2 Reasoning Model
We utilize Qwen2.5-VL-7B Bai et al. (2025) as our reasoning model . Despite Qwen2.5-VL’s superior performance in general-domain applications, direct adaptation to remote sensing imagery introduces domain knowledge discrepancy. Moreover, its reasoning capabilities require significant enhancement to address complex geospatial tasks. Therefore, we require fine-tuning it through reinforcement learning on remote sensing data.
3.3 Training
We utilize Group Relative Policy Optimization (GRPO) Shao et al. (2024) to train RemoteReasoner’s reasoning model. Unlike traditional reinforcement learning methods like PPO Schulman et al. (2017), GRPO directly compares groups of candidate outputs without an additional critic model. Given a query , GRPO samples N candidate outputs from the policy . Then it computes each output through a Reward function . To determine the relative quality of the outputs, GRPO further normalizes the obtained rewards and subsequently derives the advantage:
where denotes the advantage of relative to other sampled outputs. GRPO aims to make the model generate outputs that obtain higher advantages, then utilize them to optimize the policy :
where, and can be formulated as follows:
To adopt GRPO for our reasoning model, we design a composite reward function to optimize localization precision, object count accuracy, and output format. The total reward combines 3 critical components:
Accuracy reward measures bounding box alignment using Intersection-over-Union (IoU) with optimal Hungarian matching. For predicted boxes and ground truth :
(4) -
Count reward penalizes deviations in object quantity prediction using an exponential decay function:
(5) where and denote ground-truth and predicted counts respectively. The exponential term () creates a smooth penalty gradient that becomes stricter as relative count error increases.
Format reward checks whether RemoteReasoner’s outputs follow the specified format, returning 1 or 0 based on compliance. It is the JSON-style with reasoning paths in the think…think tag and bboxes in the answer[[,,,]…]answer tag.
Finally, the overall reward function is:
3.4 Inference Workflow
After training, the reasoning model itself generates bounding boxes (bbox) indicating target regions. To further enhance the workflow and enable both pixel-level and object-level reasoning tasks, we need to leverage established tools. Specifically, we first input the bbox and its centroid coordinates to the SAM2 Ravi et al. (2024) to generate pixel-level masks. Subsequently, morphological operations are applied to analyze the mask images, extracting the outermost polygonal vertices to obtain target contours. The mathematical formulation of this process is presented as follows:
Given a remote sensing image and a textual query , we define a multi-stage workflow for multi-task reasoning.
Geospatial Region Reasoning: A vision-language model processes the image-query pair to jointly produce:
(7) where denotes the textual reasoning trace, and is the normalized bounding box of the target region.
Geospatial Pixel Reasoning: The bounding box serves as a prompt for the Segment Anything Model (SAM) . SAM generates a binary mask isolating the target:
(8) where indicates pixel belongs to the target.
Geospatial Contour Reasoning: The target contour is derived from via morphological boundary detection:
(9) where extracts the polygonal vertices of the mask’s external boundary.
4 Experiments
| Method | ||||
| Val | Test | Val | Test | |
| SFT Methods | ||||
| LISA Lai et al. (2024) | 57.39 | 59.10 | 61.04 | 60.88 |
| PixelLM Ren et al. (2024) | 57.79 | 59.22 | 57.94 | 60.01 |
| PSALM Zhang et al. (2024) | 62.03 | 64.61 | 66.61 | 68.30 |
| SegEarth-R1 Li et al. (2025) | 64.13 | 68.25 | 68.60 | 70.75 |
| RL Method | ||||
| RemoteReasoner | 67.80 | 69.13 | 69.02 | 70.96 |
| Method | ||||
| Val | Test | Val | Test | |
| DeepSeek-VL2-tiny Guo et al. (2025) | 12.08 | 12.67 | 17.51 | 18.62 |
| GeoChat Kuckreja et al. (2024) | 10.10 | 8.89 | 12.57 | 11.44 |
| Qwen2.5-VL-7B Bai et al. (2025) | 41.21 | 45.82 | 38.77 | 41.80 |
| RemoteReasoner | 66.51 | 68.11 | 67.04 | 69.29 |
4.1 Dataset and Metrics
Dataset. We evaluate on the validation and test sets of EarthReason Li et al. (2025). Pixel-level tasks directly utilize the original annotations, while region-level and object-level tasks are transformed using M2B Liu et al. (2024a) and M2C Miao et al. (2025) annotation formats based on the original labels, respectively. Additionally, we employ the RSVG Zhan et al. (2023) test set for Visual Grounding evaluation and the RRSISD Liu et al. (2024b) test set for Referring Expression Segmentation tasks, with their training sets explicitly excluded from model training. For unseen category evaluation, we manually selected 100 images covering 10 distinct unseen categories from the RemoteSAM-270K Yao et al. (2025b) dataset.
Metrics. Following previous works Yao et al. (2025b), Li et al. (2025), we adopt , , and as region- and pixel-level evaluation metrics. For contour reasoning, we employ F1 score with 1-pixel and 3-pixel dilated kernel ( & ) as evaluation metrics, and incorporate two geometric error measures: Average Symmetric Distance () and Hausdorff Distance ().
| Method | ||||||||
| Val | Test | Val | Test | Val | Test | Val | Test | |
| EPOC Chen et al. (2024a) | 0.10 | 0.10 | 0.19 | 0.20 | 60.07 | 64.38 | 243.98 | 255.81 |
| RemoteReasoner | 0.46 | 0.44 | 0.55 | 0.58 | 41.50 | 45.20 | 160.56 | 165.11 |
4.2 Implementations Details
We employ GRPO Shao et al. (2024) to fine-tune the Qwen2.5-VL-7B-Instruct Bai et al. (2025) model using LoRA Hu et al. (2022) (rank=8, alpha=16, targeting all linear layers). Training is conducted for 24 epochs on 8 × NVIDIA L20 GPUs with a global batch size of 512 (achieved via per-device batch size 8 and gradient accumulation steps 8), using the AdamW Loshchilov and Hutter (2017) optimizer with a learning rate of 1e-6, bf16 precision, and DeepSpeed Rasley et al. (2020) Zero3 for memory efficiency.
4.3 Main Results
4.3.1 Geospatial Region Reasoning
This section validates RemoteReasoner’s performance on region-level reasoning tasks, with results presented in Tab. 2. Compared to the remote sensing multi-modal foundation model GeoChat Kuckreja et al. (2024), RemoteReasoner demonstrates at least a 56.41% improvement. This significant gain stems from GeoChat’s weak reasoning capabilities, which fail to interpret ambiguous queries. Similarly, RemoteReasoner achieves over 22.39% improvement against general-purpose MLLMs, e.g., Qwen-2.5VL-7B Bai et al. (2025) and DeepSeek-VL2-tiny (3B) Guo et al. (2025). Although these MLLMs possess basic reasoning abilities, their insufficient domain-specific knowledge in remote sensing leads to suboptimal performance on geospatial region-level reasoning tasks.
4.3.2 Geospatial Pixel Reasoning
We evaluate RemoteReasoner’s performance on pixel-level reasoning tasks using the EarthReason dataset. We compare against state-of-the-art general-domain methods (LISA Lai et al. (2024), PixelLM Ren et al. (2024), PSALM Zhang et al. (2024)) and the remote sensing-domain SegEarth-R1 Li et al. (2025), with results summarized in Tab. 1. Experimental results demonstrate that despite not directly utilizing pixel-level labels during training, our approach remarkably surpasses SFT-based methods that leverage both pixel annotations and reasoning trajectories. For instance, RemoteReasoner achieves 10.41% higher than LISA on the validation set. Compared to the best-performing SegEarth-R1, we observe a 3.67% improvement. The primary reason lies in the superior robustness of reinforcement learning (RL) frameworks compared to supervised fine-tuning (SFT) paradigms. Even without direct exposure to pixel-level annotations, our model effectively utilizes bounding boxes to prompt the Segment Anything Model (SAM), achieving precise segmentation results through this indirect yet efficient interaction mechanism.
4.3.3 Geospatial Contour Reasoning
This section validates the contour extraction performance of RemoteReasoner. We select EPOC Chen et al. (2024a) as a comparative baseline. Since EPOC lacks reasoning capabilities, we crop the bounding boxes (bbox) generated by RemoteReasoner and input them into EPOC to obtain contours, as shown in Fig. 2. Since EPOC is class-agnostic, it extracts contours for all objects within the region (e.g., first and third rows). In contrast, RemoteReasoner’s contours are derived from segmentation masks, resulting in tighter alignment with target boundaries. Furthermore, under low-resolution conditions (second row), EPOC produces blurry or disconnected contours, whereas our method maintains sharp and closed boundaries.
Quantitative results (as shown in Tab. 3) demonstrate that RemoteReasoner surpasses EPOC by at least 0.34 in both F1 scores, with significantly lower distance errors compared to EPOC. This performance gap can be attributed to EPOC’s tendency to extract extraneous contours and its failure to handle indistinct target boundaries effectively.
| Task | Method | ||
| VG | Trained | ||
| MGVLF Zhan et al. (2023) | 76.78 | 68.04 | |
| Falcon Yao et al. (2025a) | 56.87 | - | |
| SkyEyeGPT Zhan et al. (2024) | 70.50 | - | |
| LHRS-Bot Muhtar et al. (2024) | 73.45 | - | |
| Zero-shot | |||
| MiniGPT-V2 Zhu et al. (2023) | 46.64 | - | |
| Qwen-VL-Chat Hui et al. (2024) | 44.76 | - | |
| RemoteReasoner | 50.76 | 49.85 | |
| Task | Method | ||
| RES | Trained | ||
| GeoGround Zhou et al. (2024) | - | 60.50 | |
| SegEarth-R1 Li et al. (2025) | 67.56 | 66.40 | |
| Zero-shot | |||
| LISA Lai et al. (2024) | - | 26.78 | |
| PixelLM Ren et al. (2024) | - | 31.65 | |
| RemoteReasoner | 54.29 | 50.97 | |
| Category | Car | Ship | Field | Apron | Grass |
| 21.56 | 50.03 | 36.72 | 54.09 | 38.27 | |
| Category | Trucks | Hangar | Parking | Roof | Warehouses |
| 25.58 | 58.34 | 30.52 | 52.85 | 41.00 |
4.4 Further Analysis
After reinforcement learning training, can the reasoning model maintain the original capabilities of the MLLM? To explore this, we evaluate RemoteReasoner’s extrinsic capabilities, including its ability to handle unseen task types and recognize out-of-distribution categories.
4.4.1 Can RemoteReasoner Do Other Tasks?
We evaluate RemoteReasoner on Visual Grounding and Referring Expression Segmentation tasks, with results summarized in Table X. Comparative experiments are conducted against state-of-the-art LLM-based approaches, including general-domain methods (MiniGPT-V2 Zhu et al. (2023), LISA Lai et al. (2024), PixelLM Ren et al. (2024)) and remote sensing-specific frameworks (Falcon Yao et al. (2025a), LHRS-Bot Muhtar et al. (2024), GeoGround Zhou et al. (2024)). Notably, all remote sensing methods were trained on task-specific data, while RemoteReasoner’s zero-shot performance achieves 80% of their supervised results. It demonstrates our framework’s strong generalization capabilities with reinforcement learning.
4.4.2 Can RemoteReasoner Recognize Unseen Category?
Given that the EarthReason Li et al. (2025) dataset contains only 28 categories, which is significantly insufficient to cover common remote sensing targets. Therefore, we further investigate RemoteReasoner’s recognition capability for unseen categories. We manually select 100 images spanning 10 novel categories from RemoteSAM-270K Yao et al. (2025b) that are absent in EarthReason for evaluation. As represented in Tab. 5, RemoteReasoner demonstrates robust recognition of out-of-distribution categories with consistent accuracy. It demonstrates that our RL strategy avoids overfitting to category distributions in the training data. However, the model still exhibits insufficient capabilities in detecting small objects (e.g., , cars), indicating a direction for future research.
5 Conclusion
We present RemoteReasoner, a novel workflow advancing autonomous geospatial reasoning. Overcoming limitations of SFT-based approaches, it leverages RL strategy to cultivate robust reasoning capabilities while preserving MLLM generalization. Its core innovation is a unified inference pipeline enabling efficient multi-granularity task transformation (pixel-, region-, and object-level) from a single MLLM output, eliminating redundant per-task decoders. RemoteReasoner achieves SOTA results, including 40% accuracy gains in region-level reasoning, and demonstrates superior generalization across diverse, complex queries. It provides a flexible foundation for downstream geospatial intelligence requiring nuanced intent interpretation and multi-format outputs. This work is still in progress, and the model and code will be open-sourced in the next version.